OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 answers 6 −6 All real numbers No solution

OpenStudy (anonymous):

First step is to use the distributive property.

Parth (parthkohli):

Simplify, with the distributive property. \(\Large \color{MidnightBlue}{\rightarrow a(x \pm y) = ax \pm ay }\)

OpenStudy (lgbasallote):

or \[\Large \color{midnightblue}{ay \pm ax}\]

OpenStudy (anonymous):

so that means that it is 6

OpenStudy (lgbasallote):

is that what you got in your solution?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Tell me how you solved .

OpenStudy (lgbasallote):

yeah tell him..coz he's so demanding..he's le boss

Parth (parthkohli):

Meow... you there?

OpenStudy (anonymous):

yea

Parth (parthkohli):

Maybe he got it on a planet where 1 = 2

Parth (parthkohli):

How did you get the answer?

Parth (parthkohli):

Didn't choose a in a while?

OpenStudy (anonymous):

thats what mi sultion came to so is it right or not cuz i wanna kno beforei take my test on things jus like that

OpenStudy (anonymous):

\[x\ne6\]

OpenStudy (lgbasallote):

now you have 3 more choices to snipe

OpenStudy (anonymous):

then again it might be NO SOLUTION

OpenStudy (anonymous):

It might be? Where on earth did you get the notion that "it might be" is something you are allowed to say in the context of mathematics?

OpenStudy (lgbasallote):

*might*

OpenStudy (lgbasallote):

probability dude :p

OpenStudy (anonymous):

That's applied math it doesn't count igbiw

Parth (parthkohli):

From a user's approach, I'd rather flunk on a test than to cheat on it. Thanks.

OpenStudy (lgbasallote):

id rather not take the test than flunk it...that way no one will *surely* know that i didnt know the answer

Parth (parthkohli):

OR understand what it actually says.

Parth (parthkohli):

OR if I don't, then think of the probability :D

OpenStudy (lgbasallote):

yeah taking the probability is MUCH easier than algebra

Parth (parthkohli):

I can help you @kibrelle

OpenStudy (lgbasallote):

(take the blue pill)

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\rightarrow P(Event) = {\text{Preferred number of outcomes} \over \text{Total number of outcomes}} }\)

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\rightarrow P(Correct) = {1 \over 4} }\)

OpenStudy (anonymous):

i dont need help anymore with that one because it is and i know it is NO SOLUTION

Parth (parthkohli):

How did you get that answer o.O?

OpenStudy (anonymous):

How sure are you on a scale from 1 to 10?

Parth (parthkohli):

@nbouscal He isn't sure how sure he is sure that he surely got the answer correct :/

OpenStudy (lgbasallote):

how sure are you he isn't sure how sure he is sure that he surely got the answer correct?

Parth (parthkohli):

100% sure that he isn't sure the he got the answer sure correct.

OpenStudy (anonymous):

100% confidence level?! Impossibru

OpenStudy (lgbasallote):

95% of statistics are made up

OpenStudy (anonymous):

10

Parth (parthkohli):

10! Wow, you were sure, weren't you? BTW, 10...apply it to the equation.

OpenStudy (anonymous):

\[ −3(x + 3) = −3(x + 1) − 5 \\ -3x -9 = -3x -3 -5\\ -9 = -3-5 9 =8 \] So no solutions.